OpenStudy (anonymous):

a cube of glass of 5cm side and mass 306 g, has a cavity inside it. if the density of glass is 2.55gcm-3. Find the volume of the cavity.

OpenStudy (ahsome):

@asad.malik, find the volume of the glass with the cavity, and without. Then, take away each other

OpenStudy (shamim):

Volume of given cube is=5*5*5=125cm^3 Right?

OpenStudy (shamim):

Now we need a formula Mass=volume*density

OpenStudy (shamim):

Now volume of cube excluding cavity=306/2.55=?

OpenStudy (shamim):

Volume of cavity=125-306/2.55=?

OpenStudy (shamim):

Response plz!!!!